# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

### v1

@@ 3.1 Multi-Teacher Adversarial Robustness Distillation
The goal of MTARD is to learn a student with strong accuracy as the clean teacher and strong robustness as the robust teacher. In the actual operation process, however, two teachers may display different knowledge scales; Meanwhile, the student may have different learning speeds toward a different type of teacher’s knowledge. The above two points may lead to an unbalanced performance in accuracy and robustness. So how to handle multi-teacher adversarial robustness distillation becomes a problem to be solved in the following two subsections.
3.2 Entropy-Based Balance in MTARD
We can theoretically obtain a student with both accuracy and robustness by introducing two teachers. However, due to the different model architectures and training methods, the teachers have different knowledge scales. Eliminating the trade-off between accuracy and robustness will become unrealistic if the student learns from the two teachers with no equal knowledge scales. The student will enjoy simultaneously acquiring equal knowledge scales from two teachers who appear to be equally knowledgeable. Based on the above analysis, we propose the Entropy-Based Balance algorithm in MTARD.
Then how to quantify the teacher’s knowledge scale needs to be solved. Intuitively speaking, a well-trained network model is trained from a network model with randomly initialized weights. The model with randomly initialized weights can be considered as a general model $I$ without any knowledge; After training, the model can be considered to possess the knowledge brought by the different optimization processes. Based on information theory, relative entropy (also known as Kullback–Leibler divergence [30]) represents the information cost required from one distribution to another distribution. Therefore, we apply the relative entropy to denote clean teacher’s knowledge scales $K_{nat}^{T_{nat}}$ and robust teacher’s knowledge scales $K_{adv}^{T_{adv}}$ as follows:
$K_{nat}^{T_{nat}}=KL(I(x_{nat}),T_{nat}(x_{nat})),$ (6)
$K_{adv}^{T_{adv}}=KL(I(x_{adv}),T_{adv}(x_{adv})),$ (7)
where $I$ denotes the general model with random weights.
Our goal is to reduce the gap between clean teachers’ knowledge scale $K_{nat}^{T_{nat}}$ and robust teachers’ knowledge scale $K_{adv}^{T_{adv}}$, and we define the gap as $\Delta_{K}$. To further simplify the gap $\Delta_{K}$, a theorem 1 is provided, and the theorem 1 is proved in proof 1.
@@ 3.2 Entropy-Based Balance in MTARD
The knowledge scale’s difference between clean and robust teacher $\Delta_{K}$ can be expressed by the difference between the teacher’s information entropy $H(P_{nat}^{T_{nat}})$ and $H(P_{adv}^{T_{adv}})$:
$\Delta_{K}=||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (8)
Proof 1
Based on information theory, the difference in teachers’ relative entropy in knowledge distillation is as follows:
$\Delta_{K}=||K_{adv}^{T_{adv}}-K_{nat}^{T_{nat}}||.$ (9)
For a model with randomly initialized weights, the model’s prediction has no preference for any samples, so a reasonable assumption can be made: at an initial state, when the general model $I$ outputs $P_{nat}^{I}=\{p_{1}^{I}(x_{nat}),...,p_{C}^{I}(x_{nat})\}$ for $x_{nat}$ and $P_{adv}^{I}=\{p_{1}^{I}(x_{adv}),...,p_{C}^{I}(x_{adv})\}$ for $x_{adv}$, thus:
$p_{k}^{I}(x_{nat})=p_{k}^{I}(x_{adv})=1/C,k=1,2,...,C.$ (10)
Assume the clean teacher’s predicted distribution for clean examples is $P_{nat}^{T_{nat}}$, and the robust teacher’s predicted distribution for adversarial examples is $P_{adv}^{T_{adv}}$, we further expand the $\Delta_{K}$ as follows:
$\Delta_{K}$ $=||H(P_{adv}^{T_{adv}})-\sum_{k=1}^{C}p_{k}^{T_{adv}}(x_{adv})log(p_{k}^{I}(x_{adv}))$
$+\sum_{k=1}^{C}p_{k}^{T_{nat}}(x_{nat})log(p_{k}^{I}(x_{nat}))-H(P_{nat}^{T_{nat}})||,$ (11)
based on the assumption of the initial state (Eq. (10)), we have:
$\Delta_{K}=$ $||H(P_{adv}^{T_{adv}})-log(1/C)\sum_{k=1}^{C}p_{k}^{T_{adv}}(x_{adv})$
$+log(1/C)\sum_{k=1}^{C}p_{k}^{T_{nat}}(x_{nat})-H(P_{nat}^{T_{nat}})||,$ (12)
$=$ $||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (13)
Then the Theorem 1 is proved.
Based on Theorem 1, by reducing the difference in the teachers’ information entropy, teachers’ knowledge scales will tend to be consistent, which is conducive to mitigating the trade-off between accuracy and robustness. So we define the above knowledge scales gap as an optimization objective in mathematical form as follows:
$\min\Delta_{K}=||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||.$ (14)
To solve the optimization object $\min\Delta_{K}$, the intuitive strategy is to jointly re-train the teachers so that the information entropy is relatively consistent. However, the training process requires an additional calculation consumption, which is not an ideal approach.
Then the hyper-parameter temperature $\tau$ attracts our attention. To properly utilize the teacher, a common practice in the field of knowledge distillation is adjusting the temperature $\tau$, which can directly influence the teacher’s predicted distribution and produce a suitably soft label for the student. Two characteristics of the temperature make us believe that adjusting the temperature can achieve the optimization goal elegantly. Firstly, the adjustment of the temperature can directly transform the information entropy of the model’s prediction, and the effect is as shown in Fig. 2: When the temperature changes, the prediction’s information entropy also changes obviously. Secondly, the calculation cost of updating temperature is almost negligible, which means that we may achieve the final goal if a good optimization towards temperature $\tau$ is designed.
Based on the above consideration, we expect to dynamically update the multiple teachers’ temperatures in the training process to achieve our final optimization goal. More specially, our optimization goal has changed from balancing the teacher’s knowledge scales to indirectly adjusting the information entropy by adjusting the temperature.
@@
To solve the above optimization Eq. (14), we need to apply the gradient of the expected object $\Delta_{K}$ concerning the temperature $\tau_{nat}$ and $\tau_{adv}$ and update them according to the gradient descent principle in the training process. Here we apply the optimization process of temperature $\tau_{nat}$ to expand our method (the optimization process of temperature $\tau_{adv}$ is similar, and we do not repeat the description), which is formulated as follows:
@@
Here, we calculate the partial derivative of information entropy $H(P)$ ($P$ can be considered as a general expression including $P_{nat}^{T_{nat}}$ and $P_{adv}^{T_{adv}}$, and $\tau$ is the same) with respect to temperature $\tau$ and can get the result as follows:
where the $q$ represents the $exp(z(x)/\tau)$ in Eq. (3), and the more detailed derivation can be found in Appendix A (Proof A.1.). Meanwhile, we prove that the partial derivative $\nabla_{\tau}H(P)\geq 0$, and this conclusion is proved in Appendix A (Proof A.2.), thus the information entropy $H(P)$ shows a monotonically increasing relationship with the temperature $\tau$. Based on the consideration of the stable update, we use the constant 1 to estimate $\nabla_{\tau}H(P)$ in the training process of updating the teacher’s temperature, and we use $r_{\tau}$ to control the update range.
@@
On a practical level, our Entropy-Based Balance algorithm can balance the teacher’s knowledge scale by controlling the information entropy. When the information entropy of one teacher is greater or smaller than another teacher, both teachers’ temperatures will be updated until the information entropy of the two teachers reaches a relatively balanced state, and the knowledge scales of different teachers will be equally matched. With the adjustment of the Entropy-Based Balance, the clean teacher and the robust teacher can appear equally knowledgeable when imparting knowledge to the student, which is helpful to mitigate the trade-off between accuracy and robustness.
It should be mentioned that we only update the teachers’ temperatures ($\tau_{nat}$ and $\tau_{adv}$) and keep the student’s temperature ($\tau_{s}$) unchanged in the Entropy-Based Balance. In addition, in order to prevent information from being over-regulated, we set an upper and lower bound for the temperature to prevent the following situations. If the temperature is too small, the teacher’s predicted distribution will be infinitely close to the one-hot label’s distribution; When the temperature is too large, the teacher’s predicted distribution will be close to the uniform distribution.
3.3 Normalization Loss Balance in MTARD
@@
On the mathematical level, the total loss in MTARD ultimately used for the student update at time $t$ can be represented as $L_{total}(t)$, which can be formulated as follows:
@@
where the $L_{nat}(t)$ and $L_{adv}(t)$ denote the KL loss mentioned in Eq. (4), and the $w_{nat}(t)$ and $w_{adv}(t)$ denote the constant hyper-parameters $1-\alpha$ and $\alpha$ mentioned in Eq. (4) in our task if without the Normalization Loss Balance. Similar to the analysis in section 3.2, the knowledge scales’ gap between the student and the teacher can be directly reflected by the value of the clean loss $L_{nat}(t)$ and adversarial loss $L_{adv}(t)$, and the vital to control the student’s knowledge from multiple teachers consistent is controlling the loss weight of $w_{nat}(t)$ and $w_{adv}(t)$, which directly influence the student’s learning of $L_{nat}(t)$ and $L_{adv}(t)$.
@@
where $[\tilde{L}(t)]^{\beta}$ denotes the $\tilde{L}(t)$ power of $\beta$, and $\beta$ is set to control the strength of correcting the imbalance. The $r_{nat}(t)$ and $r_{adv}(t)$ strengthen the disadvantaged losses and weakens the advantaged losses to eliminate the imbalance between the student’s learning speeds in different types of knowledge. The big $\beta$ is applicable when the vibration of relative loss value is obvious, while the small $\beta$ is suitable for $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ with stable influence abilities.
@@
3.4 Overview of MTARD
All in all, to transfer both clean and robust knowledge to the student, we propose the MTARD to apply two teachers and propose the Entropy-Based Balance algorithm for teachers and the Normalization Loss Balance algorithm for the student to optimize the adversarial training process. The final minimization of MTARD is defined as follows:
@@ 3.4 Overview of MTARD
and the complete process of MTARD is in Algorithm 1.
Algorithm 1 Overview of MTARD
1:Initialize student $S$ with weight $\theta_{S}$ and temperature $\tau_{s}$, pretrained teacher $T_{nat}$ and $T_{adv}$, the initial temperature $\tau_{nat}$ and $\tau_{adv}$ for $T_{nat}$ and $T_{adv}$, respectively, the benign clean examples $x_{nat}$ and the labels $y$, the threat bound $\Omega$, the max training epochs $max$-$epoch$.
2:for $epoch=0$ to $max$-$epoch$ do
3:     for $t=0$ to $max$-$step$ do
4:         Acquire adversarial example $x_{adv}$ by Eq. (5).
5:         Get $L_{nat}(t)=KL(S(x_{nat};\tau_{s}),T_{nat}(x_{nat};\tau_{nat}))$.
6:         Get $L_{adv}(t)=KL(S(x_{adv};\tau_{s}),T_{adv}(x_{adv};\tau_{adv}))$.
7:         if $epoch=0$ and $t=0$ then
8:              Record $L_{nat}$ and $L_{adv}$ as $L_{nat}(0)$ and $L_{adv}(0)$.
9:         end if
10:         Update $w_{nat}(t)$ and $w_{adv}(t)$ by Eq. (25) and Eq. (26).
11:         Update $\theta_{S}$ by Eq. (27).
12:         Update $\tau_{nat}$ and $\tau_{adv}$ by Eq. (18) and Eq. (19).
13:     end for
14:end for

### v3

@@ 3.1 Balanced Multi-Teacher Adversarial Distillation
The goal of B-MTARD is to obtain a student with strong accuracy as the clean teacher and strong robustness as the robust teacher. In the actual operation process, however, two teachers may display different knowledge scales; Meanwhile, the student may have different learning speeds toward different types of teachers’ knowledge. The above two points may lead to an unbalanced performance in accuracy and robustness. So how to handle multi-teacher adversarial robustness distillation becomes a problem to be solved in the following two subsections.
3.2 Entropy-Based Balance in B-MTARD
3.2.1 Measuring Knowledge Scale
Due to the different model architectures and training methods, the teachers may have different knowledge scales. Eliminating the accuracy-robustness trade-off will become unrealistic if the student learns from the two teachers with no equal knowledge scales. Then we want to know how to measure the knowledge scale and further balance it.
In the knowledge distillation process, the part that teachers applied to guide students is the predicted distribution for the samples. So we can reasonably believe that the teacher’s knowledge exists in the predicted distribution for the samples, which actually represents the information understood by the teacher, e.g., the samples’ classes, the samples’ difficulty, or the relationship between the classes. In knowledge distillation framework [20], the teacher’s predicted distribution can be defined as $P=\{p_{1}(x),...,p_{C}(x)\}$, where the predicted possibility $p_{k}(x)$ of $k$-th class can be formulated as follows:
$p_{k}(x)=\frac{exp(z_{k}(x)/\tau)}{\sum_{j=1}^{C}exp(z_{j}(x)/\tau)},$ (5)
where $z_{k}(x)$ denotes the $k$-th dimension output logits of the model before the softmax layer, $\tau$ denotes the temperature applied in the training process of knowledge distillation.
Although the knowledge described above exists in the teacher’s predicted distributions, how to quantify them is not easy. Here we try to consider it from the optimization perspective. Intuitively speaking, a well-trained network model is trained from a network model with randomly initialized weights. The model with randomly initialized weights can be considered as a general model $I$ without any knowledge; After training, the model can be considered to possess the knowledge brought by the optimization processes. Based on information theory, relative entropy (also known as Kullback–Leibler divergence [36]) represents the information cost required from one distribution to another distribution. So the obtained knowledge scale can be denoted as the relative entropy from the predicted distribution of an initialized model $I$ to a well-trained model. Here we further define the teacher’s knowledge scale $K^{T}$ in mathematical form as follows:
$K^{T}=KL(I(x),T(x)),$ (6)
then we try to further simplify the knowledge scale $K^{T}$ and provide Theorem 1. And the corresponding proof can be found in the Appendix A (Proof A.1.).
@@ 3.2.1 Measuring Knowledge Scale
The teacher’s knowledge scale $K^{T}$ is negatively related to the information entropy $H(P^{T})$ of teacher’s predicted distribution and have a relationship as follows:
$K^{T}=logC-H(P^{T}),$ (7)
[…]
Theorem 1 implies that we can directly use information entropy to measure the teacher’s knowledge scales. From the perspective of information theory, information entropy is used to describe the degree of randomness and uncertainty of information, so it is reasonable to apply information entropy for quantifying the knowledge scale.
Then we further analyze the factors that affect information entropy. Based on Eq.(5), the information entropy of predicted distribution is affected by two parts, one is the output logits $z(x)$, and the other is the temperature $\tau$, both of them can change the information entropy to influence the knowledge scale. The output logits $z(x)$ is determined by the model itself, when the teacher is selected, $z(x)$ cannot be changed. While the temperature can be used as a hyper-parameter to directly change information entropy.
In order to further explore the relationship between information entropy and temperature, we calculate the partial derivative of information entropy $H(P)$ with respect to temperature $\tau$ and can get the results as follows:
where the $q$ represents the $exp(z(x)/\tau)$ in Eq.(5), and the more detailed derivation can be found in Appendix A (Proof A.2.). Meanwhile, we prove that the partial derivative $\nabla_{\tau}H(P)\geq 0$, and this conclusion is proved in Appendix A (Proof A.3.), thus the information entropy $H(P)$ shows a monotonically increasing relationship with temperature $\tau$. The effect is as shown in Fig. 2: when the temperature increases, the prediction’s information entropy also increases obviously. Then combined with Theorem 1, a larger temperature will lead to a smaller knowledge scale.
3.2.2 Balancing Knowledge Scale
After measuring the knowledge scale, we want to know the knowledge scales of different types of teachers, we conduct quantitative statistics, and the results are shown in Table I. The results denote that an obvious difference exists in the knowledge scale between the clean teacher and the robust teacher if without additional adjustment. So it is very necessary to balance the knowledge scale between different teachers. Only when two teachers appear to be equally knowledgeable, the student will equally learn abilities from both the clean teacher and the robust teacher, and finally mitigate the trade-off between accuracy and robustness.
TABLE I: The quantitative statistic of knowledge scale for clean teacher ResNet-56 and robust teacher WideResNet-34-10 on CIFAR-10. We count the teachers’ knowledge scales before and after the adjustment of the Entropy-Based balance algorithm. The results are based on the statistics of the teachers’ knowledge scales in the final training epoch.
Teacher Knowledge Scale (No Balance) Knowledge Scale (After Balance)
ResNet-56 2.2518 1.2480
WideResNet-34-10 1.3379 1.3249
Based on the above analysis, we can balance the clean teacher’s knowledge scale and robust teacher’s knowledge scale by minimizing the gap of teachers’ information entropy. Here we define the minimization of knowledge scales’ gap $\Delta_{K}$ as optimization objective in mathematical form:
$\min\Delta_{K}=||H(P_{adv}^{T_{adv}})-H(P_{nat}^{T_{nat}})||,$ (9)
where $P_{nat}^{T_{nat}}$ and $P_{adv}^{T_{adv}}$ denote the clean teacher’s predicted distribution towards $x_{nat}$ and adversarial teacher’s predicted distribution towards $x_{adv}$, respectively.
Based on the above analysis in Section 3.2.1, adjusting temperature can directly transform the information entropy of the model’s prediction with negligible calculation cost, while adjusting logits $z(x)$ needs an additional calculation consumption for re-train models. So we select to solve the above optimization goal Eq. (9) by updating the multiple teachers’ temperatures.
@@ 3.2.2 Balancing Knowledge Scale
Then we need to apply the gradient of the expected object $\Delta_{K}$ concerning the clean teacher’s temperature $\tau_{nat}$ and the robust teacher’s temperature $\tau_{adv}$ and update them according to the gradient descent principle in the training process. Here we apply the optimization process of temperature $\tau_{nat}$ to expand our method (the optimization process of temperature $\tau_{adv}$ is similar, and we do not repeat the description), which is formulated as follows:
@@
As described before (Eq.(8)), the information entropy $H(P)$ shows a monotonically increasing relationship with the temperature $\tau$. So based on the consideration of the stable update, we can reasonably use the constant 1 to estimate both the $\nabla_{\tau_{nat}}H(P_{nat}^{T_{nat}})$ and $\nabla_{\tau_{adv}}H(P_{adv}^{T_{adv}})$ in the training process of updating the teacher’s temperature, and we use $r_{\tau}$ to control the update range.
@@
On a practical level, our Entropy-Based Balance algorithm can balance the teacher’s knowledge scale by controlling the information entropy. When the information entropy of teachers is not equal, both teachers’ temperatures will be updated until the difference is eliminated: if a teacher has a high information entropy, then the teacher’s temperature will reduce to further reduce the information entropy, leading to the higher knowledge scale based on Eq.(27); on the contrary, if a teacher has a low information entropy, then the teacher’s temperature will increase to further increase the information entropy, leading to the lower knowledge scale. With the Entropy-Based Balance, the clean teacher and the robust teacher can appear equally knowledgeable when imparting knowledge to the student, which is helpful to mitigate the accuracy-robustness trade-off.
It should be mentioned that we only update the teachers’ temperatures ($\tau_{nat}$ and $\tau_{adv}$) and keep the student’s temperature ($\tau_{s}$) unchanged in the Entropy-Based Balance. In addition, in order to prevent information from being over-regulated, we set an upper and lower bound for the temperature to prevent the following situations: If the temperature is too small, the teacher’s predicted distribution will be infinitely close to the one-hot label’s distribution; When the temperature is too large, the teacher’s predicted distribution will be close to the uniform distribution.
3.3 Normalization Loss Balance in B-MTARD
@@
On the mathematical level, the total loss in B-MTARD ultimately used for the student update at time $t$ can be represented as $L_{total}(t)$, which can be formulated as follows:
@@
where the $L_{nat}(t)$ and $L_{adv}(t)$ denote the KL loss mentioned in Eq. (3), and the $w_{nat}(t)$ and $w_{adv}(t)$ denote the constant hyper-parameters $1-\alpha$ and $\alpha$ mentioned in Eq. (3) in our task without the Normalization Loss Balance. Similar to the analysis in section 3.2, the knowledge scales’ gap between the student and the teacher can be directly reflected by the value of the clean loss $L_{nat}(t)$ and adversarial loss $L_{adv}(t)$, and the vital to control the student’s knowledge from multiple teachers consistent is controlling the loss weight of $w_{nat}(t)$ and $w_{adv}(t)$, which directly influence the student’s learning of $L_{nat}(t)$ and $L_{adv}(t)$.
@@
where $[\tilde{L}(t)]^{\beta}$ denotes the $\tilde{L}(t)$ power of $\beta$, and $\beta$ is set to control the strength of correcting the imbalance. The $r_{nat}(t)$ and $r_{adv}(t)$ strengthen the disadvantaged losses and weaken the advantaged losses to eliminate the imbalance between the student’s learning speeds in different types of knowledge. The big $\beta$ is applicable when the vibration of relative loss value is obvious, while the small $\beta$ is suitable for $\tilde{L}_{nat}(t)$ and $\tilde{L}_{adv}(t)$ with stable influence abilities.
@@
3.4 Overview of B-MTARD
All in all, to transfer both clean and robust knowledge to the student, we propose the B-MTARD to apply two teachers and propose the Entropy-Based Balance algorithm for teachers and the Normalization Loss Balance algorithm for the student to optimize the Adversarial Training process. The final minimization of B-MTARD is defined as follows:
@@ 3.4 Overview of B-MTARD
and the complete process of B-MTARD is in Algorithm 1.
Algorithm 1 Overview of B-MTARD
1:Initialize student $S$ with weight $\theta_{S}$ and temperature $\tau_{s}$, pretrained teacher $T_{nat}$ and $T_{adv}$, the initial temperature $\tau_{nat}$ and $\tau_{adv}$ for $T_{nat}$ and $T_{adv}$, respectively, the benign clean examples $x_{nat}$ and the labels $y$, the threat bound $\Omega$.
2:for $t=0$ to $max$-$step$ do
3:     Acquire adversarial example $x_{adv}$ by Eq. (4).
4:     Get $L_{nat}(t)=KL(S(x_{nat};\tau_{s}),T_{nat}(x_{nat};\tau_{nat}))$.
5:     Get $L_{adv}(t)=KL(S(x_{adv};\tau_{s}),T_{adv}(x_{adv};\tau_{adv}))$.
6:     if $epoch=0$ and $t=0$ then
7:         Record $L_{nat}$ and $L_{adv}$ as $L_{nat}(0)$ and $L_{adv}(0)$.
8:     end if
9:     Update $w_{nat}(t)$ and $w_{adv}(t)$ by Eq. (19) and Eq. (20).
10:     Update $\theta_{S}$ by Eq. (21).
11:     Update $\tau_{nat}$ and $\tau_{adv}$ by Eq. (12) and Eq. (13).
12:end for
@@
TABLE II: Performance of the teacher networks used in our experiments.
Dataset Model Clean Acc FGSM PGDsat PGDtrades CW AA Type
CIFAR-10 ResNet-56 93.18% 19.23% 0% 0% 0% 0% Clean
WideResNet-34-10 84.91% 61.14% 55.30% 56.61% 53.84% 53.08% Robust
CIFAR-100 WideResNet-22-6 76.65% 4.85% 0% 0% 0% 0% Clean
WideResNet-70-16 60.96% 35.89% 33.58% 33.99% 31.05% 30.03% Robust
Tiny-ImageNet PreActResNet-34 66.05% 2.50% 0.04% $0$% Clean
PreActResNet-34 52.76% 27.05% 24.00% 24.63% 20.07% 18.92% Robust
TABLE III: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of CIFAR-10 and CIFAR-100 are based on the ResNet-18, and the results of Tiny-ImageNet are based on the PreActResNet-18. All results are the best checkpoint based on W-Robust Acc.
FGSM[14] PGDsat[31] PGDtrades[55] CW[4] AutoAttack[10]
DataSet Defense Clean Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust
CIFAR-10 Natural 94.57% 18.60% 56.59% 0% 47.29% 0% 47.29% 0% 47.29% 0% 47.29%
SAT[31] 84.20% 55.59% 69.90% 45.95% 65.08% 48.12% 66.16% 45.97% 65.09% 43.49% 63.85%
TRADES[55] 83.00% 58.35% 70.68% 52.35% 67.68% 53.83% 68.42% 50.23% 66.62% 49.03% 66.02%
ARD[13] 84.11% 58.40% 71.26% 50.93% 67.52% 52.96% 68.54% 50.15% 67.13% 48.20% 66.16%
RSLAD[61] 83.99% 60.41% 72.20% 53.94% 68.97% 55.73% 69.86% 52.67% 68.33% 50.98% 67.49%
SCORE[32] 84.43% 59.84% 72.14% 53.72% 69.08% 55.21% 69.82% 50.46% 67.45% 49.25% 66.84%
Fair-ARD[49] 83.41% 58.91% 71.16% 52.00% 67.71% 53.77% 68.59% 51.07% 67.24% 49.21% 66.31%
ABSLD[58] 83.21% 60.22% 71.72% 54.63% 68.92% 56.10% 69.66% 52.04% 67.63% 50.60% 66.91%
MTARD[59] 87.36% 61.20% 74.28% 50.73% 69.05% 53.60% 70.48% 48.57% 67.97% 46.18% 66.77%
B-MTARD (ours) 88.20% 61.42% 74.81% 51.68% 69.94% 54.40% 71.30% 49.88% 69.04% 47.44% 67.82%
CIFAR-100 Natural 75.18% 7.96% 41.57% 0% 37.59% 0% 37.59% 0% 37.59% 0% 37.59%
SAT[31] 56.16% 25.88% 41.02% 21.18% 38.67% 22.02% 39.09% 20.90% 38.53% 19.76% 37.96%
TRADES[55] 57.75% 31.36% 44.56% 28.05% 42.90% 28.88% 43.32% 24.19% 40.97% 23.26% 40.51%
ARD[13] 60.11% 33.61% 46.86% 29.40% 44.76% 30.51% 45.31% 27.56% 43.84% 25.71% 42.91%
RSLAD[61] 58.25% 34.73% 46.49% 31.19% 44.72% 32.05% 45.15% 28.21% 43.23% 26.50% 42.38%
SCORE[32] 56.40% 32.94% 44.67% 30.27% 43.34% 30.56% 43.48% 26.30% 41.35% 24.74% 40.57%
Fair-ARD[49] 57.81% 34.39% 46.10% 30.64% 44.23% 31.50% 44.66% 27.84% 42.83% 26.22% 42.02%
ABSLD[58] 56.77% 34.94% 45.86% 32.41% 44.59% 32.99% 44.88% 26.99% 41.88% 25.38% 41.08%
MTARD[59] 64.30% 31.49% 47.90% 24.95% 44.63% 26.75% 45.53% 23.42% 43.86% 21.31% 42.81%
B-MTARD (ours) 65.08% 34.21% 49.65% 28.50% 46.79% 29.94% 47.51% 25.45% 45.27% 23.98% 44.53%
Tiny-ImageNet Natural 65.03% 1.73% 33.38% 0.03% 32.53% 0.05% 32.54% 0% 32.52% 0% 32.52%
SAT[31] 50.08% 25.35% 37.72% 22.24% 36.16% 23.05% 36.57% 20.48% 35.28% 18.54% 34.31%
TRADES[55] 48.45% 23.59% 36.02% 21.59% 35.02% 22.09% 35.27% 17.33% 32.89% 16.66% 32.56%
ARD[13] 53.22% 27.97% 40.60% 24.92% 39.07% 25.71% 39.47% 21.41% 37.32% 19.91% 36.57%
RSLAD[61] 48.78% 27.26% 38.02% 25.00% 36.89% 25.45% 37.12% 20.87% 34.83% 17.11% 32.95%
SCORE[32] 10.05% 7.80% 8.93% 7.65% 8.85% 7.67% 8.86% 6.19% 8.13% 5.97% 8.01%
Fair-ARD[49] 46.64% 25.81% 36.23% 23.91% 35.28% 24.29% 35.47% 19.59% 33.12% 17.51% 32.08%
ABSLD[58] 47.21% 25.79% 36.50% 23.59% 35.40% 23.98% 35.60% 19.50% 33.36% 17.84% 32.53%
MTARD[59] 52.98% 26.41% 39.70% 22.55% 37.77% 23.41% 38.20% 19.36% 36.17% 16.84% 34.91%
B-MTARD (ours) 56.81% 28.12% 42.47% 23.93% 40.37% 24.94% 40.88% 19.69% 38.25% 17.71% 37.26%
TABLE IV: White-box robustness on CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. The results of three datasets are trained based on the MobileNet-v2. All results are the best checkpoint based on W-Robust Acc.
FGSM[14] PGDsat[31] PGDtrades[55] CW[4] AutoAttack[10]
DataSet Defense Clean Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust Robust W-Robust
CIFAR-10 Natural 93.35% 12.22% 52.79% 0% 46.68% 0% 46.68% 0% 46.68% 0% 46.68%
SAT[31] 83.87% 55.89% 69.88% 46.84% 65.36% 49.14% 66.51% 46.62% 65.25% 43.43% 63.65%
TRADES[55] 77.95% 53.75% 65.85% 49.06% 63.51% 50.27% 64.11% 46.06% 62.01% 45.18% 61.66%
ARD[13] 83.43% 57.03% 70.23% 49.50% 66.47% 51.70% 67.57% 48.96% 66.20% 46.60% 65.02%
RSLAD[61] 83.20% 59.47% 71.34% 53.25% 68.23% 54.76% 68.98% 51.78% 67.49% 50.23% 66.72%
SCORE[32] 82.32% 58.43% 70.38% 53.42% 67.87% 54.46% 68.39% 49.18% 65.75% 48.39% 65.36%
Fair-ARD[49] 82.65% 56.37% 69.51% 50.50% 66.58% 52.12% 67.39% 51.07% 66.86% 47.68% 65.17%
ABSLD[58] 82.50% 58.47% 70.49% 52.98% 67.74% 54.49% 68.50% 50.20% 66.35% 48.65% 65.58%
MTARD[59] 89.26% 57.84% 73.55% 44.16% 66.71% 47.99% 68.63% 43.42% 66.34% 41.02% 65.14%
B-MTARD (ours) 89.09% 58.79% 73.94% 47.56% 68.33% 50.44% 69.77% 46.81% 67.95% 44.58% 66.84%
CIFAR-100 Natural 74.86% 5.94% 40.40% 0% 37.43% 0% 37.43% 0% 37.43% 0% 37.43%
SAT[31] 59.19% 30.88% 45.04% 25.64% 42.42% 26.96% 43.08% 25.01% 42.10% 22.84% 41.02%
TRADES[55] 55.41% 30.28% 42.85% 23.33% 39.37% 28.42% 41.92% 27.72% 41.57% 22.61% 39.01%
ARD[13] 60.45% 32.77% 46.61% 28.69% 44.57% 29.63% 45.04% 26.55% 43.50% 24.86% 42.66%
RSLAD[61] 59.01% 33.88% 46.45% 30.19% 44.60% 31.19% 45.10% 27.98% 43.50% 26.33% 42.67%
SCORE[32] 49.38% 29.28% 39.33% 27.03% 38.21% 27.53% 38.46% 23.29% 36.34% 21.92% 35.65%
Fair-ARD[49] 59.18% 34.07% 46.63% 30.15% 44.67% 31.26% 45.22% 27.55% 43.37% 25.84% 42.51%
ABSLD[58] 56.67% 33.85% 45.26% 31.28% 43.98% 31.90% 44.29% 26.40% 41.54% 24.56% 40.62%
MTARD[59] 67.01% 32.42% 49.72% 25.14% 46.08% 27.10% 47.06% 24.14% 45.58% 21.61% 44.31%
B-MTARD (ours) 66.13% 34.36% 50.25% 28.47% 47.30% 29.82% 47.98% 26.50% 46.32% 24.01% 45.07%
Tiny-ImageNet Natural 62.09% 0.72% 31.41% 0% 31.05% 0% 31.05% 0% 31.05% 0% 31.05%
SAT[31] 49.03% 23.38% 36.21% 20.31% 34.67% 21.15% 35.09% 18.69% 33.86% 17.35% 33.19%
TRADES[55] 43.81% 20.10% 31.96% 18.16% 30.99% 18.36% 31.09% 13.47% 28.66% 12.82% 28.32%
ARD[13] 45.53% 22.88% 33.21% 20.43% 32.98% 21.00% 33.27% 16.81% 31.17% 15.40% 30.47%
RSLAD[61] 45.69% 24.09% 34.89% 22.30% 34.00% 22.74% 34.22% 18.63% 32.16% 17.09% 31.39%
SCORE[32] 28.27% 17.47% 22.87% 16.48% 22.38% 16.69% 22.48% 13.25% 20.76% 12.30% 20.29%
Fair-ARD[49] 47.24% 25.31% 36.28% 23.37% 35.31% 23.77% 35.51% 20.04% 33.64% 18.09% 32.67%
ABSLD[58] 48.08% 26.21% 37.15% 26.21% 37.15% 24.39% 36.24% 20.10% 34.09% 18.38% 33.23%
MTARD[59] 50.50% 23.94% 37.22% 20.45% 35.48% 21.20% 35.85% 17.45% 33.98% 15.06% 32.78%
B-MTARD (ours) 52.98% 25.60% 39.29% 21.58% 37.28% 22.58% 37.78% 18.08% 35.53% 15.68% 34.33%